OpenStudy (anonymous):

Kelly loves to eat chocolate pizza. Each square inch of chocolate is 30 calories. She orders a slice of chocolate pizza from a pizza with a diameter of 12 inches that is cut with a central angle of 40 degrees. How many calories does she eat? about 3,393 calories about 113 calories about 13 calories about 377 calories

OpenStudy (anonymous):

@amistre64 @ganeshie8

OpenStudy (amistre64):

the area of a circle is 1/2 the angle times the radius^2

OpenStudy (amistre64):

well, the radians of the angle that is

OpenStudy (amistre64):

when that angle is 360 (2pi) 2pi/2 * r^2 = pi r^2

OpenStudy (anonymous):

r=6

OpenStudy (amistre64):

angle/180 defines the coefficeint for k pi radians

OpenStudy (anonymous):

what do want me to do

OpenStudy (amistre64):

find that area of the slice of course

OpenStudy (amistre64):

\[\frac12\frac{\theta}{180}\pi~r^2\]

OpenStudy (amistre64):

then multiply that by 30 calories

OpenStudy (anonymous):

0.5*180/12

OpenStudy (amistre64):

\[\frac12\frac{\theta}{180}\pi~r^2\] \[\frac12\frac{40}{180}\pi~6^2\] \[\frac{20}{180}\pi~36\] \[\frac{20}{10}\pi~2\] \[2*2pi * 30~calories\]

OpenStudy (anonymous):

how do I put 2pi in the calculator

OpenStudy (blurbendy):

2 times the pi button (you should have a pi button on your calculator)

OpenStudy (amistre64):

pi is roughly 3.14159

OpenStudy (anonymous):

i got 376.99

OpenStudy (goldphenoix):

Which is very close to which answer choices?

OpenStudy (anonymous):

D the last one

OpenStudy (goldphenoix):

Correct, flexx. :) I hope amistre helped you. :)

OpenStudy (anonymous):

she did very well to

OpenStudy (goldphenoix):

Flex, give amistre the medal.

OpenStudy (goldphenoix):

I didn't really help you with anything. >_>

OpenStudy (amistre64):

ive got enough medals :)

OpenStudy (goldphenoix):

You can never have enough medals. :3

OpenStudy (anonymous):

i give who help and u my friend help too

OpenStudy (anonymous):

I deserve a medal too

OpenStudy (anonymous):

Meredith works at a horse stable. Her job is to exercise the horses. The stable has a circular area that Meredith runs the horses within. The diameter of the fencing is 16 yards. What is the length of the fence? 16π yards 4π yards 16 yards 4 yards

OpenStudy (anonymous):

thanks but i already had it answered buddy

OpenStudy (anonymous):

need help with this one

OpenStudy (anonymous):

Meredith works at a horse stable. Her job is to exercise the horses. The stable has a circular area that Meredith runs the horses within. The diameter of the fencing is 16 yards. What is the length of the fence? 16π yards 4π yards 16 yards 4 yards

OpenStudy (goldphenoix):

Maybe post this new question?

OpenStudy (anonymous):

i did